OpenStudy (anonymous):

The graph of y=2x^2-4x-2 has a y-intercept of (0,2). TRUE or FALSE ***Idk this one... please explain? thanks :)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

its FALSE??

OpenStudy (anonymous):

plug in x=0.... what's your y value?

OpenStudy (anonymous):

2(0)^2-4(0)-2 =0-0-2 =-2 y=-2 oohhh which is false!

OpenStudy (anonymous):

since it says it is 2 for y

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

ohhh thanks :)